OpenStudy (anonymous):

Enter the answer. Sylvia has a total of 16 trading cards: 8 superhero cards, 4 princess cards, 2 villain cards, and 2 wildlife cards. What is the ratio of superhero cards to the total number of trading cards in simplest form?

OpenStudy (anonymous):

fan & medal will be given

OpenStudy (babyx3boo):

1/2

OpenStudy (anonymous):

thank you

OpenStudy (babyx3boo):

Np :)

OpenStudy (anonymous):

Enter the answer A bake sale has 12 cookies, 9 pies, 8 cakes, 4 tarts, and 3 cobblers to sell. What is the ratio of the total number of baked goods to cobblers in simplest form?

OpenStudy (usukidoll):

add all of the bake goods.. what number do you get?

OpenStudy (anonymous):

12+9+8+4+3=36

OpenStudy (usukidoll):

ok so how many cobblers do you have?

OpenStudy (babyx3boo):

12

OpenStudy (usukidoll):

omg ughhh ignore that direct answer

OpenStudy (anonymous):

3 cobblers

OpenStudy (usukidoll):

yes a ratio is in the form of a:b so if we have a = 3 and b = 36 that's 3:36 we can divide this by 3 and have...

OpenStudy (usukidoll):

3/3 =? 36/3 = ?

OpenStudy (anonymous):

3/3=1 36/3=12 1:12

OpenStudy (usukidoll):

yes!

OpenStudy (anonymous):

thanks Suki(hope you don't mind me calling you that)

OpenStudy (usukidoll):

heheheh someone on roblox called me usuk usux XD

OpenStudy (anonymous):

i like Suki better

OpenStudy (usukidoll):

and i just got my 400th medal yay!